TABLES

TABLE I. The Poisson ratio defined as the ratio of lateral contraction to longitudinal expansion for uniaxial stress. Experimental values are given for comparison.

σ  σ
Predicted  Observed^{a}  Predicted  Observed^{a}
Cu  0.48  0.36  Al  0.47  0.33
Ag  0.48  0.37  Tl  0.47  0.35
Au  0.48  0.36  Sn  0.46  0.33
Mg  0.47  0.35  Pb  0.46  0.40–0.45
Zn  0.47  0.25  Pb  0.49  0.43
K  0.49  0.44